OpenStudy (anonymous):

From a shipment of 250 light-bulbs, a sample of 50 was selected at random and tested. If 10 light-bulbs in the sample were found to be defective, how many defective light-bulbs would be expected in the entire shipment?

OpenStudy (anonymous):

240

OpenStudy (anonymous):

All they did was just make a sample and check. Only 10 were bad. 240 were shipped. It's not like they said per 50 lightbubls in each sample FAN AND GIVE A MEDAL IF THIS HELPED

OpenStudy (anonymous):

That's not the right answer.

OpenStudy (anonymous):

A.30, B.40, C.75,D.50.

OpenStudy (anonymous):

Shouldn't the rate of error in the sample be proportional to the entire shipment?

OpenStudy (anonymous):

10/50 = x / 250

OpenStudy (anonymous):

I dunno... I tried solving it on my own, and got 50. 1/10= 0.5 ---> 0.5 OR 5x10= 50..?

OpenStudy (anonymous):

I tried. :/

OpenStudy (anonymous):

Well, thanks for helping!

OpenStudy (anonymous):

Good luck, お宅お兄さん

OpenStudy (anonymous):

Thanks! :D

OpenStudy (anonymous):

Maybe its 40